OpenStudy (babyslapmafro):

Please help me solve the following problem. Find the directional derivative of f at P in the direction of the vector a. f(x,y)=(y^2)ln(x) P(1,4) a=<-3,3>

OpenStudy (babyslapmafro):

This is what I have done: \[D _{a}f(x,y)=fx(x,y)(-3)+fy(x,y)(3)\] \[=\frac{ y ^{2} }{ x }(-3)+2yln(x)(3)\]

OpenStudy (babyslapmafro):

My final answer @ P: \[D _{a}f(1,4)=-48\] The answer in the book: \[-8\sqrt{2}\]

OpenStudy (babyslapmafro):

Ok I think I figured it out. I need to convert the vector (a) into a unit-vector.